Kim Distribution Services (KDS) distributes food purchased in bulk to small retailers. The firm is divided into two divisions: Purchasing and Distribution. Purchasing is responsible for ordering goods from the manufacturer, receiving them, and then moving them to the appropriate location in the warehouse. Distribution is responsible for taking orders from retailers, picking the products from the warehouse for the orders, and packaging the orders for shipment. KDS has a policy of filling every order on the day of the order. If the firm is out of a particular item, it will ship a partial order and complete the order when the item is back in stock. Occasionally, an order will not have to be packaged if the retailer chooses to take delivery at the KDS loading dock.

Distribution has 15 employees who are responsible for the activities, and all 15 are trained to handle any of the three tasks. Each of these employees works 40 hours per week for 50 weeks. There is an allowance of 15 percent of the employees’ time for training and other administrative tasks. The total costs of distribution for the coming year are estimated to be $1,055,700. When asked, the manager of Distribution estimated the following times for each of the three activities:

  • Taking orders: 10 minutes.

  • Picking orders: 14 minutes.

  • Packaging orders: 20 minutes.

During the year, the Distribution received 26,500 orders. Because of out-of-stock events, pickers had to pick 31,500 orders. 29,200 orders were packaged.

Required:
a. What is the cost per minute for activities in Distribution? (Round your answer to 2 decimal places.)

b. What is the cost of an order that requires all three activities? (Round your intermediate calculations and final answer to 2 decimal places.)

c. How many minutes of unused capacity did Distribution have for the year?

d. What was the cost of the unused capacity in Distribution? (Round your intermediate calculations to 2 decimal places and final answer to the

a. Cost per minute
b. Total costs
c. Unused minutes
d. Cost of unused capacity   
Answer #1
a. Cost per minute=Total estimated cost of distribution/Total minutes available in distribution
Total minutes available in distribution=Number of employees*Hours worked per week*Number of weeks*Minutes in an hour=15*40*50*60=1800000 minutes
Cost per minute=1055700/1800000=$ 0.59
b. Cost of an order=Total minutes required in all three activities*Cost per minute=(10+14+20)*0.59=44*0.59=$ 25.96
c. Unused minutes=Total minutes available-Total minutes used
Total minutes used:
Minutes
Tracking orders (26500*10) 265000
Picking orders (31500*14) 441000
Packaging orders (29200*20) 584000
Total 1290000
Unused minutes=1800000-1290000=510000 minutes
d. Cost of unused capacity=Unused minutes*Cost per minute=510000*0.59=$ 300900